OpenStudy (anonymous):

If a baseball player has four hits in 11 at bats, what is his batting average? 0.400 0.364 0.357 0.345 Javier earned the following scores on his first five tests: 77, 89, 84, 80, and 85. What score did he earn on his last test to have a test mean of 84%? 89% 84% 86% 88% A set of data can have no mode. true or false. What is the range of the data below? 45, 19, 23, 67, 28, 35, 46, 21, 58, 60, 23, 51 44 48 41 37 Which data set has a median of 15? 9, 17, 13, 15, 16, 8, 12 18, 15, 11, 14, 19, 15, 6 7, 16, 14, 16, 11, 7, 17 18, 9, 19, 16, 6, 20. 13.

OpenStudy (anonymous):

didnt get to say this in my first box but im not ery good at math and im trying to get my grade up so please help.

OpenStudy (anonymous):

*very

OpenStudy (yanasidlinskiy):

A range is the difference between the smallest number and the biggest number which is what?

OpenStudy (anonymous):

what you mean which is what?

OpenStudy (yanasidlinskiy):

What is the difference between the biggest number and the smallest #?

OpenStudy (anonymous):

6

OpenStudy (yanasidlinskiy):

Ok. So that would be your answer for the range! What do YOU think. Can a set of data have no mode??

OpenStudy (anonymous):

umm 6 cant be the anwser its not a choice on my question and im pretty sure it has to have a mode.

OpenStudy (yanasidlinskiy):

It doesn't have to have a mode. So for your A set of data can have no mode.- true. And what are your answer choices!?

OpenStudy (anonymous):

you can see them its the vertical line of numbers under every question.

OpenStudy (yanasidlinskiy):

Ok. back to the range. Did you subtract the bigger# from the smallest #?? I'm sorry I didn't correct you in the first place.

OpenStudy (anonymous):

yes 51 - 45 right

OpenStudy (yanasidlinskiy):

No. 51 is NOT the biggest #. Look again. And 45 is NOT the smallest #. Look again on that one too.

OpenStudy (anonymous):

oh oops thought it just got bigger one side so i didnt look in the middle

OpenStudy (anonymous):

60-19=41

OpenStudy (yanasidlinskiy):

Ok. you are correct!!!. So now to the median. Do you know what a median is??

OpenStudy (anonymous):

no

OpenStudy (yanasidlinskiy):

Set the data #'s from least to greatest. When thinking of median think of middle. Once you set the data from least to greatest, what is the number that is in the middle??

OpenStudy (yanasidlinskiy):

For the median, you can rule out automatically c and d. Because they don't have 15 at all. So now you are left with a and B. Which one is it??

OpenStudy (anonymous):

so would it be like 15div by 2= 7 and a half

OpenStudy (texaschic101):

1. 11/4 = 0.3636 rounds to 0.364 2. (77 + 89 + 84 + 80 + 85 + x)/6 = 84 (415 + x)/6 = 84 --- multiply by 6 415 + x = 504 x = 504 - 415 x = 89 % 3. True...a set of data does not have to have a mode 4. To get the range of a set of data, subtract the smallest number from the largest number. 67 - 19 = 48 5. First put your numbers in numerical order. The median is the middle number.....if your data has an odd number of numbers, then the median will be in the middle. If the data set has an even number of numbers, then there will be 2 middle numbers...you then add them and divide by 2. Your data set has an odd number of numbers, therefore, the set that has 15 in the center is the correct answer.

OpenStudy (anonymous):

can you just show me what you are typing

OpenStudy (anonymous):

ok thank you I was wondering if you were really there

OpenStudy (texaschic101):

no....your data set has an odd number of numbers. There is no need to divide. Just find the middle number. Your data set has 7 numbers.....that means 3 numbers on each side of the middle number....so basically the 4th number is your median

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

so 15

OpenStudy (texaschic101):

The only time you would have to divide by 2 to find your median, is if your data set has 2 middle numbers....the only way there will be 2 middle numbers, is if there is an even number of numbers in your data set